an annuity pays 6.55% annual interest rate compunded semiannually. If $500 is deposited every six months for two years, how much interest is earned?
Answer by stanbon(48535)

an annuity pays 6.55% annual interest rate compunded semiannually. If $500 is deposited every six months for two years, how much interest is earned?
------------------
A(t) = P(1+(r/n))^(nt)
A(2) = 500(1+(0.0655/2))^(2*2)
A(2) = 500(1.1376)
This is the value of the account after 2 years.
-------------------------
Interest = 0.1376*500 = $68.79
====================================
Cheers,
Stan H.